What is the frequency of a photon with a wavelength of 1040 nm. Report your answers to three significant digits.

The frequency is = _________ s-1
1
Expert's answer
2019-11-12T12:01:46-0500

Wavelength λ =1040nm= 1.04*10^-6

"C = \u03bbf" ; where C is the speed of light =3*10^8.

Frequency of photon"f = C\/ \u03bb"

 f = 3*10^8/1.04*10^-6

f= 2.89*10^14 Hz